OpenStudy (anonymous):

Evaluate –1x + 7y + 4x + 1y for x = 4 and y = –2.

OpenStudy (anonymous):

-4-14+16-2=-4

Directrix (directrix):

3(4) + 8(-2)= 12 - 16 = -4